# Interact with me: Joint Egocentric Forecasting of Intent to Interact, Attitude and Social Actions

### v1

@@ Abstract
For efficient human-agent interaction, an agent should proactively recognize their target user and prepare for upcoming interactions. We formulate this challenging problem as the novel task of jointly forecasting a person’s intent to interact with the agent, their attitude towards the agent and the action they will perform, from the agent’s (egocentric) perspective. So we propose SocialEgoNet - a graph-based spatiotemporal framework that exploits task dependencies through a hierarchical multitask learning approach. SocialEgoNet uses whole-body skeletons (keypoints from face, hands and body) extracted from only 1 second of video input for high inference speed. For evaluation, we augment an existing egocentric human-agent interaction dataset with new class labels and bounding box annotations. Extensive experiments on this augmented dataset, named JPL-Social, demonstrate real-time inference and superior performance (average accuracy across all tasks: 83.15%) of our model outperforming several competitive baselines. The additional annotations and code will be available upon acceptance.
@@ II Related Work
Intent prediction. Modeling and predicting human intent based on visual input is critical for social agents to effectively serve or collaborate with humans [8]. In such tasks, nonverbal cues such as body pose, head movements, and facial expressions are crucial. Studies often focus on modeling the temporal evolution of these cues to understand the dynamic aspects of human intent [9]. Multimodal data, such as gaze and motion cues have been employed to analyse object manipulation intent [10]. Gaze and pose are also used to predict user engagement using a support vector classifier with a social robot [11]. As for modeling the temporal evolution of the non-verbal cues, a Hidden Semi-Markov Model was used to reason about block placement intentions [12], while LSTM has been used to predict users intent to interact with a service robot [13]. In general, recurrent models are widely used and shown to capture the temporal patterns in non-verbal cues effectively.
Action prediction. While action prediction is a well-studied area in computer vision, our focus is specifically on action prediction from an egocentric perspective. This task typically involves anticipating hand activities [14, 15] and employs either LSTM or transformer-based approaches. LSTM-based methods often use rolling-unrolling techniques [16], but they struggle to model long-horizon temporal dependencies. This limitation has been addressed by goal-based learning [17] and memory-based modules [18]. Transformers are increasingly being used for action prediction [15] due to their superior performance and the ease of integrating multiple modalities, especially text. The integration of large language models (LLMs) has been explored in several studies [19, 20], highlighting the potential for enhanced multimodal understanding in action anticipation.
@@ III JPL-Social: Augmented JPL Dataset
We use a pre-trained YOLOv3 [21] to detect all visible subjects in every frame of a video. The tracks thus created (one for every person) are further annotated. The video-level labels provided originally are assigned to the interacting person directly. For the subjects not interacting, their tracks are annotated with two new actions: gaze (i.e., looking at the observer) and no response (i.e., focusing on their own business and ignoring the observer). Next, we annotate the intent and attitude of each subject. The intent of the interacting subject is labelled as ‘interacting’; for others we consider two labels: interested and not interested. The interested label is assigned to subjects that are not yet interacting but are interested in interacting in the near future. A person is considered interested when constantly looking at the observer (camera), otherwise, the person is labelled as not interested. The not interested label is assigned subjects that completely ignore the observer. We create another set of labels that indicates the attitude of a subject towards the observer. This is done based on the actions they perform and applicable to only the ’interacting’ subjects. Punching or Throwing an object to the observer are labelled as having a negative attitude while the rest are considered as having a positive attitude (see Fig. 3). Overall, we have 290 person-level tracks from all visible subjects in 200 video clips that are annotated with three labels each: their intent to interact (3 classes), their attitude toward the observer (2 classes), and the (inter)action performed (10 classes).
@@ 8       Pass the combined feature $\boldsymbol{z}^{(t)}$ to MSA
TABLE II: Comparison of SocialEgoNet with competitive spatiotemporal models on the JPL-Social dataset by observing the first second of interaction (30 frames). The best results are in bold, while the second-best results are underlined. “$\Delta$” refers to the time to extract the whole-body pose keypoints. AlphaPose [22] takes 1.27 ms to extract whole-body pose features from an annotated frame.

### v2

@@ Abstract
For efficient human-agent interaction, an agent should proactively recognize their target user and prepare for upcoming interactions. We formulate this challenging problem as a novel task of jointly forecasting a person’s intent to interact with the agent, their attitude towards the agent and the action they will perform, from the agent’s (egocentric) perspective. We propose SocialEgoNet - a graph-based spatiotemporal framework that exploits task dependencies through a hierarchical multitask learning approach. SocialEgoNet uses whole-body skeletons (keypoints from face, hands and body) extracted from only 1 second of video input for high inference speed. For evaluation, we augment an existing egocentric human-agent interaction dataset with new class labels and bounding box annotations. Extensive experiments on this augmented dataset, named JPL-Social, demonstrate real-time inference and superior performance (average accuracy across all tasks: 83.15%) of our model outperforming several competitive baselines. The additional annotations and code are available at github.com/biantongfei/SocialEgoNet.
@@ II Related Work
Intent and attitude prediction. Modeling and predicting human intent based on visual input is critical for social agents to effectively serve or collaborate with humans [8]. In such tasks, nonverbal cues such as body pose, head movements, and facial expressions are crucial. Studies often focus on modeling the temporal evolution of these cues to understand the dynamic aspects of human intent [9]. Cues, such as gaze and motion have been used to analyse object manipulation intent [10]. Gaze and pose are also used to predict user engagement using a support vector classifier with a social robot [11]. As for modeling the temporal evolution of the non-verbal cues, a Hidden Semi-Markov Model was used to reason about block placement intentions [12], while LSTM has been used to predict users intent to interact with a service robot [13]. In general, recurrent models are widely used and shown to capture the temporal patterns in non-verbal cues effectively. Facial expressions and body gestures convey valuable information to perceive social intents and attitudes [14, 15]. CNNs have demonstrated strong performance on facial feature extraction [16]. Also, CNNs can integrate facial and body pose features, with LSTM effectively capturing temporal dynamics for user attitude perception [17].
Action prediction. While action prediction is a well-studied area in computer vision, our focus is specifically on action prediction from an egocentric perspective. This task typically involves anticipating hand activities [18, 19] and employs either LSTM or transformer-based approaches [20]. LSTM-based methods often use rolling-unrolling techniques [21], but they struggle to model long-horizon temporal dependencies. This limitation has been addressed by goal-based learning [22] and memory-based modules [23]. Transformers are increasingly being used for action prediction [19] due to their superior performance and the ease of integrating multiple modalities, especially text. The integration of large language models (LLMs) has been explored in several studies [24, 25], highlighting the potential for enhanced multimodal understanding in action anticipation.
@@ III JPL-Social: Augmented JPL Dataset
We use a pre-trained YOLOv3 [26] to detect all visible subjects in every frame of a video. The tracks thus created (one for every person) are further annotated. The video-level labels provided originally are assigned to the interacting person directly. For the subjects not interacting, their tracks are annotated with two new actions: gaze (i.e., looking at the observer) and no response (i.e., focusing on their own business and ignoring the observer). Next, we annotate the intent and attitude of each subject. The intent of the interacting subject is labelled as ‘interacting’; for others we consider two labels: interested and not interested. The interested label is assigned to subjects that are not yet interacting but are interested in interacting in the near future. A person is considered interested when constantly looking at the observer (camera), otherwise, the person is labelled as not interested. The not interested label is assigned to subjects that completely ignore the observer. We create another set of labels that indicates the attitude of a subject towards the observer. This is done based on the actions they perform and applicable to only the ’interacting’ subjects. Punching or Throwing an object to the observer are labelled as having a negative attitude while the rest are considered as having a positive attitude (see Fig. 3). Overall, we have 290 person-level tracks from all visible subjects in 200 video clips that are annotated with three labels each: their intent to interact (3 classes), their attitude toward the observer (2 classes), and the (inter)action performed (10 classes).
@@ 8       Pass the combined feature $\boldsymbol{z}^{(t)}$ to MSA
TABLE II: Comparison of SocialEgoNet with competitive spatiotemporal models on the JPL-Social dataset by observing the first second of interaction (30 frames). The best results are in bold, while the second-best results are underlined. “$\Delta$” refers to the time to extract the whole-body pose keypoints. AlphaPose [27] takes 1.27 ms to extract whole-body pose features from an annotated frame. All the experiments run on a RTX 4070.